# Accuracy of the Ensemble Kalman Filter in the Near-Linear Setting

### v1

@@ 2.1 The Algorithm
The ensemble Kalman filter may be derived as a particle approximation of various mean field dynamics [calvello2025ensemble]. The specific mean field ensemble Kalman filter that we study in this paper reads
@@
where $\widehat{\pi}^{\rm EK}_{j+1}={\rm Law}(\widehat{u}_{j+1},\widehat{y}_{j+1})$ and independence of the noise terms (1.5) is still assumed to hold, as in (1.5). See Subsection 1.2 for the definition of the covariance matrices that appear in (2.1c). We denote by $\mu^{\rm EK}_{j}$ the law of $u_{j}$. In order to rewrite the evolution of $\mu_{j}^{\rm EK}$ in terms of maps on probability measures we introduce, for a given $y^{\dagger}_{j+1}$, the map $\mathsf{T}_{j}\colon\mathcal{P}^{2}_{\succ 0}(^{d_{u}}\times^{d_{y}})\to\mathcal{P}(^{d_{u}})$ defined by
@@
Here the subscript denotes the pushforward and, for given $\pi\in\mathcal{P}(^{d_{u}}\times^{d_{y}})$, $z\in^{d_{y}}$,
$\mathscr{T}(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet},\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};\pi,z)\colon$${}^{d_{u}}\times^{d_{y}}\to^{d_{u}};$
$(u,y)\mapsto$ $\,\,u+\mathcal{C}^{uy}(\pi)\mathcal{C}^{yy}(\pi)^{-1}\bigl{(}z-y\bigr{)}.$ (2.3)
Evolution of the probability measure $\mu_{j}^{\rm EK}$ may then be written as [calvello2025ensemble]
@@
We now discuss the preceding map in relation to (1.11). The specific affine map $\mathscr{T}$ used in (2.1c) is determined by measure $\widehat{\pi}^{\rm EK}_{j+1}$ (here equal to $\mathsf{Q}\mathsf{P}\mu_{j}^{\rm EK}$) and data $y^{\dagger}_{j+1}$. The following observations explain why the law of $u_{j}$ in (2.1) evolves according to (2.4):
  • if $u_{j}\sim\mu_{j}^{\rm EK}$, then $\widehat{u}_{j+1}\sim\mathsf{P}\mu_{j}^{\rm EK}$, by definition of $\mathsf{P}$;
  • random vector $(\widehat{u}_{j+1},\widehat{y}_{j+1})$ is distributed according to $\widehat{\pi}^{\rm EK}_{j+1}=\mathsf{Q}\mathsf{P}\mu_{j}^{\rm EK}$;
  • equation (2.1c) then implies that $u_{j+1}\sim\mathsf{T}_{j}\mathsf{Q}\mathsf{P}\mu_{j}^{\rm EK}$.
As we show in Lemma B.7, operator $\mathsf{T}_{j}$ coincides with conditioning operator $\mathsf{B}_{j}$ over the Gaussians $\mathcal{G}(^{d_{u}}\times^{d_{y}})\subset\mathcal{P}(^{d_{u}}\times^{d_{y}}).$ In the particular case where $\mu_{0}$ is Gaussian, which is a standing assumption in this paper, and the operators $\Psi$ and $h$ are linear, the mean field ensemble Kalman filter (2.4) thus reproduces the exact filtering distribution (1.11), given by the Kalman filter itself. The aim of this paper is to analyze the accuracy of ensemble Kalman methods when $\Psi$ and $h$ are not assumed to be linear.
The ensemble Kalman filter as implemented in practice may be derived as a particle approximation of the mean field dynamics as defined by sample paths in (2.1) or as an evolution on measures in (2.4). For any $\pi$ in the range of $\mathsf{Q}$, $\mathcal{C}^{yy}(\pi)=\mathcal{C}^{hh}(\pi)+\Gamma$ and $\mathcal{C}^{uy}(\pi)=\mathcal{C}^{uh}(\pi)$, using the notation introduced in, and following, (1.3). This is since the noise in the observation component defining $\pi$ is then independent of the state component defining $\pi.$ Using this the particle approximation of (2.1) takes the form
@@
Our theorems in Subsections 2.2 and 2.3 concern relationships between the true filter (1.11) and the mean field ensemble Kalman filter (2.4). We study the setting in which $\Psi$ and $h$ are not assumed to be linear, so that the true filter is not Gaussian, in Subsection 2.2; and then we study small perturbations away from the Gaussian setting that arise when the vector fields $\Psi$ and $h$ are close to affine in Subsection 2.3. The theorem in Subsection 2.4 concerns the relationship between the true filter (1.11) and the ensemble Kalman filter (2.6). In this subsection, we combine existing analysis on the convergence of the ensemble Kalman filter to the mean field ensemble Kalman filter with the results from the previous subsections to state and prove an error estimate between the ensemble Kalman filter and the true filter in the non-linear setting. Specifically, we study the case of a vector field $\Psi$ that is a bounded perturbation away from affine and an affine vector field $h$. To state our theorems we will use the following set of assumptions:
@@
There exists positive constants $\kappa_{y},\kappa_{\Psi},\kappa_{h},\ell_{h},\sigma$ and $\gamma$ such that the data $\{y_{j}^{\dagger}\}$, the vector fields $(\Psi,h)$ and the covariances $(\Sigma,\Gamma)$ satisfy:
@@ 2.2 Stability Theorem: Mean Field Ensemble Kalman Filter
Roughly speaking, our stability theorem states that if the true filtering distributions $(\mu_{j})_{j\in\llbracket 0,J-1\rrbracket}$ are close to Gaussian, after appropriate lifting to the state/data space, then the distribution $\mu^{\rm EK}_{j}$ given by the mean field ensemble Kalman filter (2.4) is close to the true filtering distribution $\mu_{j}$ given by (1.11) for all $j\in\llbracket 0,J\rrbracket$.
@@
Assume that the probability measures $(\mu_{j})_{j\in\llbracket 0,J\rrbracket}$ and $(\mu^{\rm EK}_{j})_{j\in\llbracket 0,J\rrbracket}$ are obtained respectively from the dynamical systems (1.11) and (2.4), initialized at the same Gaussian probability measure $\mu_{0}=\mu_{0}^{\rm EK}\in~{}\mathcal{G}(^{d_{u}})$. That is,
@@
    For any probability measure $\mu$ with finite first and second order polynomial moments $\mathcal{M}_{1}(\mu)$ and $\mathcal{M}_{2}(\mu)$, the means of the probability measures $\mathsf{P}\mu$ and $\mathsf{Q}\mathsf{P}\mu$ are bounded from above, and their covariances are bounded both from above and from below. The constants in these bounds depend only on the parameters $\kappa_{\Psi}$, $\kappa_{h}$, $\Sigma,\Gamma$ and on $\mathcal{M}_{1}(\mu)$ and $\mathcal{M}_{2}(\mu)$. See Lemmas B.1 and B.2.
@@
    Let $(\mu_{j})_{j\in\llbracket 1,J\rrbracket}$ and $(\mu^{\rm EK}_{j})_{j\in\llbracket 1,J\rrbracket}$ denote the probability measures obtained respectively from the dynamical systems (1.11) and (2.4), initialized at the same Gaussian measure $\mu_{0}=\mu_{0}^{\rm EK}\in~{}\mathcal{G}(^{d_{u}})$. Then for any $q\in\mathbb{N}$, there exist constants $M_{q},M_{q}^{\rm EK}<\infty$ depending on $\mathcal{M}_{q}(\mu_{0}),\kappa_{y},\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma,J$ so that
@@
    For any Gaussian measure $\mu\in\mathcal{G}(^{d_{u}}\times^{d_{y}})$, it holds that $\mathsf{B}_{j}\mu=\mathsf{T}_{j}\mu$. See Lemma B.7 and also [calvello2025ensemble].
@@
    The map $\mathsf{P}$ is Lipschitz on $\mathcal{P}(^{d_{u}})$ for the metric $d_{g}$, with a Lipschitz constant $L_{\mathsf{P}}$ depending on the parameters $\kappa_{\Psi}$ and $\Sigma$. See Lemma B.8.
@@
    The map $\mathsf{Q}$ is Lipschitz on $\mathcal{P}(^{d_{u}})$ for the metric $d_{g}$, with a Lipschitz constant $L_{\mathsf{Q}}$ depending on the parameters $\kappa_{h}$ and $\Gamma$. See Lemma B.9.
@@
    The map $\mathsf{B}_{j}$ satisfies, for any $\pi\in\{\mathsf{Q}\mathsf{P}\mu:\mu\in\mathcal{P}(^{d_{u}})\,\text{and }\mathcal{M}_{2}(\mu)<\infty\}\subset\mathcal{P}(^{d_{u}}\times^{d_{y}})$, the bound
@@
    where $C_{\mathsf{B}}=C_{\mathsf{B}}\bigl{(}\mathcal{M}_{2}(\mu),\kappa_{y},\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma\bigr{)}$. This statement concerns the stability of the $\mathsf{B}_{j}$ operator between a measure and its Gaussian approximation. See Lemma B.10.
@@
    The map $\mathsf{T}_{j}$ satisfies the following bound: for all $R\geqslant 1$, it holds for all probability measures $\pi\in\mathcal{P}_{R}(^{d_{u}}\times^{d_{y}})$ and $p\in\{\mathsf{Q}\mathsf{P}\mu:\mu\in\mathcal{P}(^{d_{u}})\,\text{and }\mathcal{M}_{\max\{3+d_{u},4+d_{y}\}}(\mu)<\infty\}\subset\mathcal{P}(^{d_{u}}\times^{d_{y}})$ that
@@
    for a constant $L_{\mathsf{T}}=L_{\mathsf{T}}\bigl{(}R,\mathcal{M}_{\max\{3+d_{u},4+d_{y}\}}(\mu),\kappa_{y},\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma\bigr{)}$. This statement may be viewed as a local Lipschitz continuity result in the case where the second argument of $d_{g}$ is restricted to the range of $\mathsf{Q}\mathsf{P}$. See Lemma B.11.
@@
In what follows we refer to the preceding itemized list to clarify the proof. For notational simplicity it is helpful to define the following measure of the difference between the true filtering distribution and its Gaussian projection:
@@
Assume throughout the following that $j\in\llbracket 0,J-1\rrbracket.$ The proof rests on the following use of the triangle inequality:
@@
We have used the fact that $\mathsf{T}_{j}\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j}=\mathsf{B}_{j}\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j}$ by Item 3 (Lemma B.7). Item 2 (Lemmas B.3 and B.5) shows that, for any $q\in\mathbb{N}$, there exist constants $M_{q},M_{q}^{\rm EK}<\infty$ depending on $\mathcal{M}_{q}(\mu_{0}),\kappa_{y},\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma,J$ so that
@@
Therefore by Item 1 (Lemma B.2), there is a constant $R\geqslant 1$, depending on the covariance matrices $\Sigma$, $\Gamma$, the bounds $\kappa_{\Psi}$ and $\kappa_{h}$ from Assumption H, and the moment bounds $M_{2}$ and $M_{2}^{\rm EK}$, such that for any $j\in\llbracket 0,J-1\rrbracket$ it holds that $\mathsf{Q}\mathsf{P}\mu_{j},\mathsf{Q}\mathsf{P}\mu_{j}^{\rm EK}\in\mathcal{P}_{R}(^{d_{u}}\times^{d_{y}})$. By Items 4, 5 and 7 (Lemmas B.11, B.9 and B.8), the first term in (2.8a) satisfies
$d_{g}\left(\mathsf{T}_{j}\mathsf{Q}\mathsf{P}\mu_{j}^{\rm EK},\mathsf{T}_{j}\mathsf{Q}\mathsf{P}\mu_{j}\right)$ $\leqslant L_{\mathsf{T}}\bigl{(}R,M_{\max\{3+d_{u},4+d_{y}\}}\bigr{)}L_{\mathsf{Q}}L_{\mathsf{P}}d_{g}\left(\mu_{j}^{\rm EK},\mu_{j}\right),$ (2.9)
where, for conciseness, we omitted the dependence of the constants on $\kappa_{y},\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma$.
Since it holds that $\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j}\in\mathcal{G}_{R}(^{d_{u}}\times^{d_{y}})$ by definition of $R$, the second term in (2.8a) may be bounded using Item 7 (Lemma B.11) and the definition in (2.7) of $\varepsilon$:
$d_{g}\left(\mathsf{T}_{j}\mathsf{Q}\mathsf{P}\mu_{j},\mathsf{T}_{j}\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j}\right)\leqslant L_{\mathsf{T}}\bigl{(}R,M_{\max\{3+d_{u},4+d_{y}\}}\bigr{)}d_{g}\left(\mathsf{Q}\mathsf{P}\mu_{j},\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j}\right)\leqslant L_{\mathsf{T}}\bigl{(}R,M_{\max\{3+d_{u},4+d_{y}\}}\bigr{)}\varepsilon.$
Finally, the third term in (2.8a) can be bounded using item 6 (Lemma B.10) and the definition in (2.7) of $\varepsilon$:
@@
Therefore, letting $\ell=L_{\mathsf{T}}\bigl{(}R,M_{\max\{3+d_{u},4+d_{y}\}}\bigr{)}L_{\mathsf{Q}}L_{\mathsf{P}}$, we have shown that
$d_{g}(\mu^{\rm EK}_{j+1},\mu_{j+1})$ $\leqslant\ell d_{g}(\mu^{\rm EK}_{j},\mu_{j})+\bigl{(}L_{\mathsf{T}}\bigl{(}R,M_{\max\{3+d_{u},4+d_{y}\}}\bigr{)}+C_{\mathsf{B}}\bigr{)}\varepsilon$
and the conclusion follows from the discrete Grönwall lemma, using the fact that $\mu_{0}^{\rm EK}=\mu_{0}.$

### v2

@@ 2.1 The Algorithm
The ensemble Kalman filter as implemented in practice may be derived as a particle approximation of various mean field dynamics [calvello2025ensemble]. The particular mean field ensemble Kalman filter that we consider in this paper may be written as
@@
where $\widehat{\pi}^{{\rm EK}}_{j+1}={\rm Law}(\widehat{u}_{j+1},\widehat{y}_{j+1})$ and where the independence condition (1.5) holds. We refer the reader back to Subsection 1.2 for the definition of the covariance matrices appearing in (2.1c). We denote by $\mu^{{\rm EK}}_{j}$ the law of $u_{j}$. We aim to formulate the evolution of $\mu_{j}^{{\rm EK}}$ in terms of maps on probability measures, hence we introduce, for a given $y^{\dagger}_{j+1}$, the map $\mathsf{T}_{j}\colon\mathcal{P}^{2}_{\succ 0}(^{d_{u}}\times^{d_{y}})\to\mathcal{P}(^{d_{u}})$ defined by
@@
Here for given $\pi\in{\mathcal{P}^{2}_{\succ 0}}(^{d_{u}}\times^{d_{y}})$, $z\in^{d_{y}}$, the transport map $\mathscr{T}$ is defined as
$\mathscr{T}(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet},\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};\pi,z)\colon$${}^{d_{u}}\times^{d_{y}}\to^{d_{u}};$ (2.3a)
$(u,y)\mapsto$ $\,\,u+\mathcal{C}^{uy}(\pi)\mathcal{C}^{yy}(\pi)^{-1}\bigl{(}z-y\bigr{)}.$ (2.3b)
Evolution of the probability measure $\mu_{j}^{{\rm EK}}$ may then be written (see [calvello2025ensemble]) as
@@
We note that a specific instance of map $\mathscr{T}$ defined in (2.3) appears in (2.1c) and is determined by probability measure $\pi:=\widehat{\pi}^{{\rm EK}}_{j+1}$ (here equal to $\mathsf{Q}\mathsf{P}\mu_{j}^{{\rm EK}}$) and data $z:=y^{\dagger}_{j+1}$. We refer to [carrillo2022ensemble] for a step-by-step argument detailing why the law of $u_{j}$ in (2.1) evolves according to (2.4). The operator $\mathsf{T}_{j}$ is equivalent, over the Gaussians $\mathcal{G}(^{d_{u}}\times^{d_{y}})\subset\mathcal{P}(^{d_{u}}\times^{d_{y}}),$ to the conditioning operator $\mathsf{B}_{j}$; see Lemma B.7. We recall that in the particular case where $\mu_{0}$ is Gaussian, which we assume throughout the paper, and the operators $\Psi$ and $h$ are linear, the mean field ensemble Kalman filter (2.4) coincides with the filtering distribution (1.9). However, this paper focuses on the aim of analyzing the accuracy of ensemble Kalman methods when $\Psi$ and $h$ are not assumed to be linear.
The ensemble Kalman filter as implemented in practice may be derived as a particle approximation of the mean field dynamics as defined by sample paths in (2.1) or as an evolution on measures in (2.4). To this end, we observe that for any $\pi$ in the range of $\mathsf{Q}$, $\mathcal{C}^{yy}(\pi)=\mathcal{C}^{hh}(\pi)+\Gamma$ and $\mathcal{C}^{uy}(\pi)=\mathcal{C}^{uh}(\pi)$, using the notation introduced in, and following, (1.3). This is since the noise in the observation component defining $\pi$ is then independent of the state component defining $\pi.$ Using this observation, the particle approximation of (2.1) takes the form
@@
Our theorems in Subsections 2.2 and 2.3 regard the relationship between the true filter (1.9) and the mean field ensemble Kalman filter (2.4). In Subsection 2.2 we consider the setting in which $\Psi$ and $h$ exhibit linear growth but are not assumed to be linear, hence the true filter is not Gaussian; in Subsection 2.3 we then consider small perturbations of the Gaussian setting that arise when the vector fields $\Psi$ and $h$ are close to affine. The theorem in Subsection 2.4 concerns the relationship between the true filter (1.9) and the ensemble Kalman filter (2.6). In this subsection, we combine existing analysis on the convergence of the ensemble Kalman filter to the mean field ensemble Kalman filter with the results from the previous subsections to state and prove an error estimate between the ensemble Kalman filter and the true filter in the nonlinear setting. Specifically, we study the case of a vector field $\Psi$ that is a bounded perturbation away from affine and an affine vector field $h$. To state our theorems we will use the following set of assumptions.
@@
There exist positive constants $\kappa_{y},\kappa_{\Psi},\kappa_{h},\ell_{h},\sigma$ and $\gamma$ such that the data $\{y_{j}^{\dagger}\}$, the vector fields $(\Psi,h)$ and the covariances $(\Sigma,\Gamma)$ satisfy:
@@ 2.2 Stability Theorem: Mean Field Ensemble Kalman Filter
In this section we consider the setting of dynamics and observation operator satisfying the linear growth assumptions contained in Assumption H. Informally the result of Theorem 2.1 shows that, under these assumptions, and if the true filtering distributions $(\mu_{j})_{j\in\llbracket 0,J-1\rrbracket}$ are close to Gaussian after lifting to the joint state and data space, then the distribution $\mu^{{\rm EK}}_{j}$ given by the mean field ensemble Kalman filter (2.4) is close to the filtering distribution $\mu_{j}$ given by (1.9) for all $j\in\llbracket 0,J\rrbracket$.
@@
Assume that the probability measures $(\mu_{j})_{j\in\llbracket 0,J\rrbracket}$ and $(\mu^{{\rm EK}}_{j})_{j\in\llbracket 0,J\rrbracket}$ are given respectively by the dynamical systems (1.9) and (2.4), initialized at the Gaussian probability measure $\mu_{0}=\mu_{0}^{{\rm EK}}\in~{}\mathcal{G}(^{d_{u}})$. That is,
@@
    For any probability measure $\mu$ with finite first and second order polynomial moments $\mathcal{M}_{1}(\mu)$ and $\mathcal{M}_{2}(\mu)$, the means of $\mathsf{P}\mu$ and $\mathsf{Q}\mathsf{P}\mu$ are bounded from above, and their covariances are bounded both from above and from below. The constants in these bounds depend only on the parameters $\kappa_{\Psi}$, $\kappa_{h}$, $\Sigma,\Gamma$ and on $\mathcal{M}_{1}(\mu)$ and $\mathcal{M}_{2}(\mu)$. See Lemmas B.1 and B.2.
@@
    Let $(\mu_{j})_{j\in\llbracket 1,J\rrbracket}$ and $(\mu^{{\rm EK}}_{j})_{j\in\llbracket 1,J\rrbracket}$ denote the probability measures obtained respectively from the dynamical systems (1.9) and (2.4), initialized at the same Gaussian measure $\mu_{0}=\mu_{0}^{{\rm EK}}\in~{}\mathcal{G}(^{d_{u}})$. Then for any integer $q\geqslant 2$, there exist constants $M_{q},M_{q}^{{\rm EK}}<\infty$ depending on $\mathcal{M}_{q}(\mu_{0}),\kappa_{y},\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma,J$ so that
@@
    For any Gaussian measure $\mu\in\mathcal{G}(^{d_{u}}\times^{d_{y}})$, we have that $\mathsf{B}_{j}\mu=\mathsf{T}_{j}\mu$. See Lemma B.7 and [calvello2025ensemble, carrillo2022ensemble].
@@
    The map $\mathsf{P}$ is Lipschitz on $\mathcal{P}(^{d_{u}})$ for the metric $d_{g}$. The Lipschitz constant $L_{\mathsf{P}}$ depends only on the parameters $\kappa_{\Psi}$ and $\Sigma$. See Lemma B.8.
@@
    The map $\mathsf{Q}$ is Lipschitz on $\mathcal{P}(^{d_{u}})$ for the metric $d_{g}$. The Lipschitz constant $L_{\mathsf{Q}}$ depends only on the parameters $\kappa_{h}$ and $\Gamma$. See Lemma B.9.
@@
    The map $\mathsf{B}_{j}$ satisfies that, for all $\pi\in\{\mathsf{Q}\mathsf{P}\mu:\mu\in\mathcal{P}(^{d_{u}})\,\text{and }\mathcal{M}_{2}(\mu)<\infty\}\subset\mathcal{P}(^{d_{u}}\times^{d_{y}})$, it holds
@@
    where $C_{\mathsf{B}}=C_{\mathsf{B}}\bigl{(}\mathcal{M}_{2}(\mu),\kappa_{y},\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma\bigr{)}$. Namely, this item regards the stability of the operator $\mathsf{B}_{j}$ between a probability measure and its Gaussian projection. See Lemma B.10.
@@
    The map $\mathsf{T}_{j}$ satisfies the following bound: for all $R\geqslant 1$, it holds for all $\pi\in\mathcal{P}_{R}(^{d_{u}}\times^{d_{y}})$ and $p\in\{\mathsf{Q}\mathsf{P}\mu:\mu\in\mathcal{P}(^{d_{u}})\,\text{and }\mathcal{M}_{2\cdot\max\{3+d_{u},4+d_{y}\}}(\mu)<\infty\}\subset\mathcal{P}(^{d_{u}}\times^{d_{y}})$ that
@@
    for a constant $L_{\mathsf{T}}=L_{\mathsf{T}}\bigl{(}R,\mathcal{M}_{2\cdot\max\{3+d_{u},4+d_{y}\}}(\mu),\kappa_{y},\kappa_{\Psi},\kappa_{h},{\ell_{h}},\Sigma,\Gamma\bigr{)}$. Namely, this item may be regarded as a local Lipschitz continuity result. See Lemma B.11.
@@
Within the following argument we make reference to the list of items outlined above. We begin by defining for ease of exposition the following measure of difference between the true filter and its Gaussian projection:
@@
We assume throughout the argument that $j\in\llbracket 0,J-1\rrbracket.$ The proof rests upon the following application of the triangle inequality:
@@
Here we have used the fact that $\mathsf{T}_{j}\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j}=\mathsf{B}_{j}\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j}$ by Item 3 (Lemma B.7). Item 2 (Lemmas B.3 and B.5) shows that, for any integer $q\geqslant 2$, there exist constants $M_{q},M_{q}^{{\rm EK}}<\infty$ depending on $\mathcal{M}_{q}(\mu_{0}),\kappa_{y},\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma,J$ so that
@@
Therefore by Item 1 (Lemma B.2), there is a constant $R\geqslant 1$, depending on the covariances $\Sigma$, $\Gamma$, the bounds $\kappa_{\Psi}$ and $\kappa_{h}$ from Assumption H, and the moment bounds $M_{2}$ and $M_{2}^{{\rm EK}}$, such that for any $j\in\llbracket 0,J-1\rrbracket$ it holds that $\mathsf{Q}\mathsf{P}\mu_{j},\mathsf{Q}\mathsf{P}\mu_{j}^{{\rm EK}}\in\mathcal{P}_{R}(^{d_{u}}\times^{d_{y}})$. In view of Items 4, 5 and 7 (Lemmas B.11, B.9 and B.8), the first term in (2.8a) satisfies
$d_{g}\left(\mathsf{T}_{j}\mathsf{Q}\mathsf{P}\mu_{j}^{{\rm EK}},\mathsf{T}_{j}\mathsf{Q}\mathsf{P}\mu_{j}\right)$ $\leqslant L_{\mathsf{T}}\bigl{(}R,M_{2\cdot\max\{3+d_{u},4+d_{y}\}}\bigr{)}L_{\mathsf{Q}}L_{\mathsf{P}}d_{g}\left(\mu_{j}^{{\rm EK}},\mu_{j}\right),$ (2.9)
where, for conciseness, we have omitted the dependence of the constants in the bound on $\kappa_{y},\kappa_{\Psi},\kappa_{h},\Sigma,\Gamma$. By definition of $R$ it holds that $\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j}\in\mathcal{G}_{R}(^{d_{u}}\times^{d_{y}})$, hence the second term in (2.8a) may be bounded using Item 7 (Lemma B.11) and the definition in of $\varepsilon$ in (2.7). Indeed, it holds that
$d_{g}\left(\mathsf{T}_{j}\mathsf{Q}\mathsf{P}\mu_{j},\mathsf{T}_{j}\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j}\right)\leqslant L_{\mathsf{T}}\bigl{(}R,M_{2\cdot\max\{3+d_{u},4+d_{y}\}}\bigr{)}d_{g}\left(\mathsf{Q}\mathsf{P}\mu_{j},\mathsf{G}\mathsf{Q}\mathsf{P}\mu_{j}\right)\leqslant L_{\mathsf{T}}\bigl{(}R,M_{2\cdot\max\{3+d_{u},4+d_{y}\}}\bigr{)}\varepsilon.$
Using Item 6 (Lemma B.10) and the definition of $\varepsilon$ in (2.7) we establish the following bound on the third term in (2.8a):
@@
Therefore, setting $\ell=L_{\mathsf{T}}\bigl{(}R,M_{2\cdot\max\{3+d_{u},4+d_{y}\}}\bigr{)}L_{\mathsf{Q}}L_{\mathsf{P}}$, we have that
$d_{g}(\mu^{{\rm EK}}_{j+1},\mu_{j+1})$ $\leqslant\ell d_{g}(\mu^{{\rm EK}}_{j},\mu_{j})+\bigl{(}L_{\mathsf{T}}\bigl{(}R,M_{\max\{3+d_{u},4+d_{y}\}}\bigr{)}+C_{\mathsf{B}}\bigr{)}\varepsilon.$
The conclusion then follows from the discrete Grönwall lemma, using the fact that $\mu_{0}^{{\rm EK}}=\mu_{0}.$